OpenStudy (anonymous):

If log(7x+5)/log(2)=1 , then x=

OpenStudy (zzr0ck3r):

\[\frac{\log(7x+5)}{\log(2)}=1\\\log(7x+5)=\log(2)\] now raise both sides to be powers of what ever the base is.

OpenStudy (zzr0ck3r):

you with me @kmoses24 ?

OpenStudy (anonymous):

I know I am @zzr0ck3r, continue please :P

OpenStudy (zzr0ck3r):

what is the base of the log?

OpenStudy (anonymous):

10 @zzr0ck3r